# Topological phases of unitary dynamics: Classification in Clifford category

### v2

@@ 1. Introduction
which is likely not surjective. Indeed, if we believe in Walker’s conjecture [21] that any strictly local commuting Hamiltonian in two dimensions whose ground state subspace is locally indistinguishable can only realize a Levin–Wen model [22], then the QCA in [18] does not belong to the image of (2). Also, the QCA in [19] does not belong to the image of (2) assuming a different conjecture. On the other hand, the same conjecture by Walker implies that the map (2) is injective for ${\mathsf{d}}=3$.
@@
We present a complete classification of translation invariant Clifford QCA in all dimensions for any prime $p$, modulo a stable equivalence relation by translation invariant Clifford circuits of finite depth and shifts where the circuits and shifts are allowed to obey a coarser translation invariance. That is, we say two TI Clifford QCA are equivalent $\alpha\simeq\beta$ if $(\alpha\otimes I_{{q^{\prime}}\mathbb{Z}^{{\mathsf{d}}}})(\beta\otimes I_{{q^{\prime\prime}}\mathbb{Z}^{{\mathsf{d}}}})^{-1}$ is a Clifford circuit combined with a shift, that is overall translation invariant under a subgroup of ${\mathbb{Z}}^{\mathsf{d}}$ of finite index. The set of all equivalence classes is denoted by $\mathfrak{C}({\mathsf{d}},p)$. More formally, given a prime $p$, let ${{q}\mathbb{Z}^{{\mathsf{d}}}}$ denote the ${\mathsf{d}}$-dimensional lattice of qudits $({\mathbb{C}}^{p})^{\otimes q}$. For any translation invariant Clifford QCA $\alpha$ on ${{q}\mathbb{Z}^{{\mathsf{d}}}}$ and a lattice ${{q^{\prime}}\mathbb{Z}^{{\mathsf{d}}}}$ we have a stabilization: if $q\leq q^{\prime}$, then $\alpha\mapsto\alpha\otimes I$ on ${{q}\mathbb{Z}^{{\mathsf{d}}}}\sqcup{{(q^{\prime}-q)}\mathbb{Z}^{{\mathsf{d}}}}={{q^{\prime}}\mathbb{Z}^{{\mathsf{d}}}}$. Modulo Clifford circuits and shifts with respect to the unaltered translation group ${\mathbb{Z}}^{\mathsf{d}}$, we have the direct limit ${\mathfrak{U}}^{-}(R)$ under the stabilization where $R={\mathbb{F}}_{p}[x_{1},\tfrac{1}{x_{1}},\ldots,x_{\mathsf{d}},\tfrac{1}{x_{\mathsf{d}}}]$.444 We have not explained what ${\mathfrak{U}}^{-}(R)$ is, but it will become clear in §3. Literally, it is the quotient group of all TI Clifford QCA on the infinite space modulo TI Clifford circuits of finite depth and shifts with stabilization. On top of that, if $q^{\prime}=b^{\mathsf{d}}q$ for some integer $b\geq 1$, then ${{q^{\prime}}\mathbb{Z}^{{\mathsf{d}}}}$ is a lattice with a supersite ${\mathbb{C}}^{p^{q}}\times\{1,2,\ldots,b\}^{\mathsf{d}}$ sitting at every point of ${\mathbb{Z}}^{d}$. Then we have a coarse-graining map ${\phi_{\#}^{({b})}}:\alpha\mapsto\alpha^{\prime}$ where $\alpha^{\prime}$ is the same as $\alpha$ but with the supersite interpreted as a new site. This coarse-graning makes a directed system of stable groups ${\mathfrak{U}}^{-}(R)$ via ${\phi_{\#}^{({\cdot})}}$ with respect to the partial order of integers $b$ given by division. Let $\mathfrak{C}({\mathsf{d}},p)$ be the direct limit. Our main result is555 One may find it not general enough that we take translation subgroups of a special form; however, any subgroup of ${\mathbb{Z}}^{\mathsf{d}}$ of finite index is given by an integer ${\mathsf{d}}$-by-${\mathsf{d}}$-matrix of nonzero determinant, and further coarse-graining using the adjugate matrix shows that our choice of ${\phi_{\#}^{({b})}}$ is sufficient.
@@
Acknowledgments. I thank Mike Freedman, Matt Hastings, and Zhenghan Wang for useful discussions. Data availability does not apply to this research since no data was used or generated.
@@
Given a free $R$-module $R^{m}=\bigoplus_{j=1}^{m}Re_{j}$ we define a dual $(R^{m})^{*}$, the set of all $R$-linear functionals on $R^{m}$, with a standard basis $f_{1},\ldots,f_{m}:R^{m}\to R$ defined by $f_{j}(e_{j})=1$ but $f_{j}(e_{i})=0$ if $j\neq i$. The ring action on the dual is given by $(r\cdot f)(\circ)=f(\bar{r}\,\circ)$.777 In addition to this ring action, there is another ring action which is $(rf)(\circ)=rf(\circ)$. If $R$ is noncommutative, the two actions are clearly distinguished. Suppose we have a right module $M$; since we write a module element as a column matrix, this is appropriate. Then, the dual $M^{*}$ is naturally a left module. With the involution that is an antiautmorphism of $R$, we make the dual into a right module: if $f\in M^{*}$, then for $r\in R$, we define $(f\cdot r)(\circ)=f(\circ\,\bar{r})$. This is our ring action on the dual. Since our $R$ is commutative, there is no distinction between left and right modules, but it continues to be true that there are two different actions on a dual. We always use the one by the involution. Since the dual $(R^{m})^{*}$ is a free $R$-module, we may represent its elements by a column matrix using the standard basis. When doing so, one needs to be careful since elements of $R^{m}$ are also represented by column matrices. If a column matrix $v$ represents an element of $(R^{m})^{*}$, then the linear functional by $v$ is $R^{m}\ni w\mapsto v^{\dagger}w\in R$. We see that the ring action on $(R^{m})^{*}$ is consistent: $(av)^{\dagger}w=\bar{a}v^{\dagger}w$ for any $a\in R$.
@@
$\overline{\cdots}$ ${\mathbb{F}}_{p}$-linear involution of the base ring by which $x_{i}\mapsto\tfrac{1}{x_{i}}$
@@
Direct sum. When applied to concrete unitary matrices,
upper left blocks are collected in the upper left.
@@
Direct sum. When applied to matrices of forms,
the summands are placed block diagonally as usual.
$\sim$ Equivalence of quadratic $\mp$-forms by adding even forms $\mu\pm\mu^{\dagger}$
$\cong$ Congruence without stabilization
$\simeq$ Witt equivalence. Stable congruence including $\sim$
@@
Unitary groups. Tilde for quadratic, $\mathsf{E}$ for elementary.
@@
${\mathfrak{U}}^{\mp},{\widetilde{\mathfrak{U}}}^{\mp}$ Unitary groups modulo elementary ones. Tilde for quadratic.
@@ 2. Quadratic and hermitian forms
We have to find an inverse for every element $[\phi]\in{\widetilde{\mathfrak{W}}}^{\mp}(R)$. Let $\Delta=\phi\mp\phi^{\dagger}$. Let $\psi$ be a quadratic form such that $\psi\sim\Delta^{-1}\phi\Delta^{-1}$. It follows that $\psi\mp\psi^{\dagger}=\Delta^{-1}(\phi\mp\phi^{\dagger})\Delta^{-1}=\Delta^{-1}$. Note that $\psi^{\dagger}\phi\psi\sim\psi^{\dagger}(\Delta\psi\Delta)\psi$ and $\psi\phi\psi^{\dagger}\sim\psi(\Delta\psi\Delta)\psi^{\dagger}$, but $(\psi\mp\psi^{\dagger})\Delta\psi=\psi=\psi\Phi(\psi\mp\psi^{\dagger})$, which implies that $\psi^{\dagger}\Delta\psi=\psi\Delta\psi^{\dagger}$. So, $\psi\phi\psi^{\dagger}\sim\psi\Delta\psi\Delta\psi^{\dagger}=\psi\Delta\psi^{\dagger}\Delta\psi=\psi^{\dagger}\Delta\psi\Delta\psi\sim\psi^{\dagger}\phi\psi$. Using this equivalence, we see that
@@
For a quadratic form $\phi$ on a projective $R$-module $Q$, a direct summand $A$ of $Q$ such that there exists some sesquilinear function $\theta$ with which $(\phi+\theta\pm\theta^{\dagger})(a,a)=0$ for all $a\in A$, is called a sublagrangian. If $\phi$ is nonsingular, a lagrangian is a sublagrangian such that the orthogonal complement is itself.999 It is important that a lagrangian is a direct summand. Since every exact “stabilizer module” [31] has its orthogonal complement equal to itself, if we do not demand that a lagrangian be a direct summand, every exact stabilizer module would be a lagrangian. The phenomenology of exact stabilizer modules is much richer than that of direct summand lagrangians. In a sense, the classification of Clifford QCA in this paper is the classification of direct summand lagrangians over $R$. We similarly define sublagrangians and lagrangians with respect to a nonsingular hermitian forms by the same condition where the even orthogonality is replaced by just orthogonality. Note that, unlike in the classical theory of quadratic forms over fields where a quadratic form is defined to be a certain function on a vector space, our quadratic form restricted to a sublagrangian is not identically zero; it is merely required to become zero on some equivalent ($\sim$) quadratic form.
@@
Since $L$ is a direct sumamnd there is a direct complement $N$ such that $Q=N\oplus L$. Then,
@@
This means that two self-orthogonal submodules, jointly generaing the full module, are lagrangians with respect to a hermitian form. Under the conditions of this lemma, we will say that $Y$ and $Z$ are a nonsingular pair with respect to $\lambda$.
@@
(12) $\lambda=\begin{Bmatrix}Y^{*}\\ Z^{*}\end{Bmatrix}\begin{pmatrix}0&\beta\\ \gamma&0\end{pmatrix}\begin{Bmatrix}Y\\ Z\end{Bmatrix}$
@@ 3. Unitary groups and Clifford circuits
${\mathcal{H}}_{\mp}=\begin{pmatrix}0&\hat{1}^{-{\dagger}}\\ \mp\hat{1}&0\end{pmatrix}$ $\in{\mathsf{U}}^{\mp}(1)\subseteq{\mathsf{U}}^{\mp}(q)\text{ where }\alpha\in{\mathsf{GL}}(q;R)$
@@
Let us very briefly review the transcription of Pauli groups to $R$-modules [31]. The group of all finitely supported Pauli operator on ${\mathbb{Z}}^{\mathsf{d}}$ with $q$ qudits ${\mathbb{C}}^{p}$ per site is nonabelian, but its commutator subgroup consists of $p$-th root of unity. Since the single-site Pauli group abelianizes to ${\mathbb{F}}_{p}^{2q}$, the abelianization of the Pauli group on the infinite lattice is the additive group of $R^{2q}$. The translation group action gives an $R$-module structure to $R^{2q}$ that matches the usual ring action on the free $R$-module $R^{2q}$. It is our convention that upper half block of a Laurent polynomial column matrix, (i.e., $R^{q}\oplus 0\subseteq R^{q}\oplus R^{q}$), corresponds to the “$X$”-part, operator factors $X=\sum_{j\in{\mathbb{F}}_{p}}\ket{j+1}\bra{j}$ | and its powers, and the lower half block $0\oplus R^{q}$ to the “$Z$”-part, operator factors $Z=\sum_{j\in{\mathbb{F}}_{p}}e^{2\pi\mathrm{i}j/p}\ket{j}\bra{j}$ |. For example,
@@
If a ${\mathbb{C}}$-algebra automorphism $\beta$ maps a Pauli operator to a Pauli operator, we obtain a map $U$ from a Laurent polynomial column matrix to another. If $\beta$ is translation invariant, then the induced map $U$ must be $R$-linear so we have an $R$-module automorphism on $R^{2q}$. The size of a Pauli operator under $\beta$ can only be finite, and since the translation invariant map $\beta$ is determined by finitely may images, the locality blowup is uniformly bounded across the whole lattice ${\mathbb{Z}}^{\mathsf{d}}$. Therefore, $\beta$ is a Clifford QCA. Since any ${\mathbb{C}}$-algebra automorphism must preserve commutation relation, the induced $R$-automorphism $U$ must preserve $\lambda$. We hence obtain a map from Clifford QCA to $\lambda$-unitaries. Introducing auxiliary qudits in the lattice on which a QCA acts by identity amounts to our stabilization.
@@
It suffices to consider $U^{\prime}=I$ because we may multiply $(U^{\prime})^{-1}$ everywhere. By definition of ${\mathsf{U}}^{\mp}(q)$, we must have $U=\begin{pmatrix}I&Y\\ 0&I\end{pmatrix}$ where $Y^{\dagger}\mp Y=0$. If $U\in{\widetilde{\mathsf{U}}}^{\mp}(q)$, then $Y=\theta\pm\theta^{\dagger}$ for some $\theta$. ∎
@@
Suppose $L\subset Q\oplus Q^{*}$ is a lagrangian of a trivial quadratic $\mp$-form $\eta$ on $Q\oplus Q^{*}$ where $Q$ is a finitely generated projective $R$-module. Then, there exists an $\eta^{\mp}$-unitary $U$ on $Q\oplus Q^{*}$ such that $U(Q\oplus 0)=L$ and such $U$ is unique up to ${\widetilde{\mathsf{EU}}}^{\mp}(Q)$. Similarly, if $L$ is self-orthogonal with respect to the trivial form $\lambda^{\mp}$ on $Q\oplus Q^{*}$, then there exists a $\lambda^{\mp}$-unitary $U$ such that $U(Q\oplus 0)=L$, unique up to ${\mathsf{EU}}^{\mp}(Q)$.

### v3

@@ 1. Introduction
which is likely not surjective. Indeed, if we believe in Walker’s conjecture [21] that any strictly local commuting Hamiltonian in two dimensions whose ground state subspace is locally indistinguishable can only realize a Levin–Wen model [22], then the QCA in [18] does not belong to the image of (2). Also, the QCA in [19] does not belong to the image of (2) assuming a different conjecture that the $3+1$d topological order of emergent fermionic charges and fermionic loops cannot be realized in any $3$d lattice Hamiltonian [19, 20]. On the other hand, Walker’s conjecture implies that the map (2) is injective for ${\mathsf{d}}=3$.
@@
We present a complete classification of translation invariant Clifford QCA in all dimensions for any prime $p$, modulo a stable equivalence relation by translation invariant Clifford circuits of finite depth and shifts where the circuits and shifts are allowed to obey a coarser translation invariance. Given a prime $p$, let ${{q}\mathbb{Z}^{{\mathsf{d}}}}$ denote the ${\mathsf{d}}$-dimensional lattice of qudits $({\mathbb{C}}^{p})^{\otimes q}$. We say two TI Clifford QCA are equivalent $\alpha\simeq\beta$ if $(\alpha\otimes I_{{q^{\prime}}\mathbb{Z}^{{\mathsf{d}}}})(\beta\otimes I_{{q^{\prime\prime}}\mathbb{Z}^{{\mathsf{d}}}})^{-1}$ is a Clifford circuit combined with a shift, that is overall translation invariant under a subgroup of ${\mathbb{Z}}^{\mathsf{d}}$ of finite index. The set of all equivalence classes is denoted by $\mathfrak{C}({\mathsf{d}},p)$. For any translation invariant Clifford QCA $\alpha$ on ${{q}\mathbb{Z}^{{\mathsf{d}}}}$ and a lattice ${{q^{\prime}}\mathbb{Z}^{{\mathsf{d}}}}$ we have a stabilization: if $q\leq q^{\prime}$, then $\alpha\mapsto\alpha\otimes I$ on ${{q}\mathbb{Z}^{{\mathsf{d}}}}\sqcup{{(q^{\prime}-q)}\mathbb{Z}^{{\mathsf{d}}}}={{q^{\prime}}\mathbb{Z}^{{\mathsf{d}}}}$. Modulo Clifford circuits and shifts with respect to the unaltered translation group ${\mathbb{Z}}^{\mathsf{d}}$, we have the direct limit ${\mathfrak{U}}^{-}(R)$ under the stabilization where $R={\mathbb{F}}_{p}[x_{1},\tfrac{1}{x_{1}},\ldots,x_{\mathsf{d}},\tfrac{1}{x_{\mathsf{d}}}]$.444 We have not explained what ${\mathfrak{U}}^{-}(R)$ is, but it will become clear in §3. Literally, it is the quotient group of all TI Clifford QCA on the infinite space modulo TI Clifford circuits of finite depth and shifts with stabilization. On top of that, if $q^{\prime}=b^{\mathsf{d}}q$ for some integer $b\geq 1$, then ${{q^{\prime}}\mathbb{Z}^{{\mathsf{d}}}}$ is a lattice with a supersite ${\mathbb{C}}^{p^{q}}\times\{1,2,\ldots,b\}^{\mathsf{d}}$ sitting at every point of ${\mathbb{Z}}^{d}$. Then we have a coarse-graining map ${\phi_{\#}^{({b})}}:\alpha\mapsto\alpha^{\prime}$ where $\alpha^{\prime}$ is the same as $\alpha$ but with the supersite interpreted as a new site. This coarse-graning makes a directed system of stable groups ${\mathfrak{U}}^{-}(R)$ via ${\phi_{\#}^{({\cdot})}}$ with respect to the partial order of integers $b$ given by division. Let $\mathfrak{C}({\mathsf{d}},p)$ be the direct limit. Our main result is555 One may find it not general enough that we take translation subgroups of a special form; however, any subgroup of ${\mathbb{Z}}^{\mathsf{d}}$ of finite index is given by an integer ${\mathsf{d}}$-by-${\mathsf{d}}$-matrix of nonzero determinant, and further coarse-graining using the adjugate matrix shows that our choice of ${\phi_{\#}^{({b})}}$ is sufficient.
@@
Acknowledgments. I thank Mike Freedman, Matt Hastings, and Zhenghan Wang for useful discussions.
@@
Given a free $R$-module $R^{m}=\bigoplus_{j=1}^{m}Re_{j}$ we define a dual $(R^{m})^{*}$, the set of all $R$-linear functionals on $R^{m}$, with a standard basis $f_{1},\ldots,f_{m}:R^{m}\to R$ defined by $f_{j}(e_{j})=1$ but $f_{j}(e_{i})=0$ if $j\neq i$. The ring action on the dual is given by $(r\cdot f)(\circ)=f(\bar{r}\,\circ)$.777 In addition to this ring action, there is another ring action which is $(rf)(\circ)=rf(\circ)$. If $R$ is noncommutative, the two actions are clearly distinguished. Suppose we have a right module $M$; since we write a module element as a column matrix, this is appropriate. Then, the dual $M^{*}$ is naturally a left module. With the involution that is an antiautomorphism of $R$, we make the dual into a right module: if $f\in M^{*}$, then for $r\in R$, we define $(f\cdot r)(\circ)=f(\circ\,\bar{r})$. This is our ring action on the dual. Since our $R$ is commutative, there is no distinction between left and right modules, but it continues to be true that there are two different actions on a dual. We always use the one by the involution. Since the dual $(R^{m})^{*}$ is a free $R$-module, we may represent its elements by a column matrix using the standard basis. When doing so, one needs to be careful since elements of $R^{m}$ are also represented by column matrices. If a column matrix $v$ represents an element of $(R^{m})^{*}$, then the linear functional by $v$ is $R^{m}\ni w\mapsto v^{\dagger}w\in R$. We see that the ring action on $(R^{m})^{*}$ is consistent: $(av)^{\dagger}w=\bar{a}v^{\dagger}w$ for any $a\in R$.
@@
$\overline{\cdots}$ ${\mathbb{F}}_{p}$-linear involution of the base ring by which $x_{i}\mapsto\tfrac{1}{x_{i}}$.
@@
Direct sum (when applied to concrete unitary matrices,
upper left blocks are collected in the upper left).
@@
Direct sum (when applied to matrices of forms,
the summands are placed block diagonally as usual).
$\sim$ Equivalence of quadratic $\mp$-forms by adding even forms $\mu\pm\mu^{\dagger}$.
$\cong$ Congruence without stabilization.
$\simeq$ Witt equivalence; stable congruence including $\sim$.
@@
Unitary groups; tilde for quadratic, $\mathsf{E}$ for elementary.
@@
${\mathfrak{U}}^{\mp},{\widetilde{\mathfrak{U}}}^{\mp}$ Unitary groups modulo elementary ones; tilde for quadratic.
@@ 2. Quadratic and hermitian forms
We have to find an inverse for every element $[\phi]\in{\widetilde{\mathfrak{W}}}^{\mp}(R)$. Let $\Delta=\phi\mp\phi^{\dagger}$. Let $\psi$ be a quadratic form such that $\psi\sim\Delta^{-1}\phi\Delta^{-1}$. It follows that $\psi\mp\psi^{\dagger}=\Delta^{-1}(\phi\mp\phi^{\dagger})\Delta^{-1}=\Delta^{-1}$. Note that $\psi^{\dagger}\phi\psi\sim\psi^{\dagger}(\Delta\psi\Delta)\psi$ and $\psi\phi\psi^{\dagger}\sim\psi(\Delta\psi\Delta)\psi^{\dagger}$, but $(\psi\mp\psi^{\dagger})\Delta\psi=\psi=\psi\Delta(\psi\mp\psi^{\dagger})$, which implies that $\psi^{\dagger}\Delta\psi=\psi\Delta\psi^{\dagger}$. So, $\psi\phi\psi^{\dagger}\sim\psi\Delta\psi\Delta\psi^{\dagger}=\psi\Delta\psi^{\dagger}\Delta\psi=\psi^{\dagger}\Delta\psi\Delta\psi\sim\psi^{\dagger}\phi\psi$. Using this equivalence, we see that
@@
For a quadratic form $\phi$ on a projective $R$-module $Q$, a direct summand $A$ of $Q$ such that there exists some sesquilinear function $\theta$ with which $(\phi+\theta\pm\theta^{\dagger})(a,a)=0$ for all $a\in A$, is called a sublagrangian. If $\phi$ is nonsingular, a lagrangian is a sublagrangian such that the orthogonal complement is itself.999 It is important that a lagrangian is a direct summand. Since every exact “stabilizer module” [31] has its orthogonal complement equal to itself, if we do not demand that a lagrangian be a direct summand, every exact stabilizer module would be a lagrangian. The phenomenology of exact stabilizer modules is much richer than that of direct summand lagrangians. In a sense, the classification of Clifford QCA in this paper is the classification of direct summand lagrangians over $R$. We similarly define sublagrangians and lagrangians with respect to nonsingular hermitian forms by the same condition where the even orthogonality is replaced by just orthogonality. Note that, unlike in the classical theory of quadratic forms over fields where a quadratic form is defined to be a certain function on a vector space, our quadratic form restricted to a sublagrangian is not identically zero; it is merely required to become zero on some equivalent ($\sim$) quadratic form.
@@
Since $L$ is a direct summand there is a direct complement $N$ such that $Q=N\oplus L$. Then,
@@
This means that two self-orthogonal submodules with respect to a hermitian form, jointly generating the full module, are lagrangians with respect to the hermitian form. Under the conditions of this lemma, we will say that $Y$ and $Z$ are a nonsingular pair with respect to $\lambda$.
@@
(12) $\lambda=\begin{Bmatrix}X^{*}\\ Y^{*}\end{Bmatrix}\begin{pmatrix}0&\beta\\ \gamma&0\end{pmatrix}\begin{Bmatrix}X\\ Y\end{Bmatrix}$
@@ 3. Unitary groups and Clifford circuits
${\mathcal{H}}_{\mp}=\begin{pmatrix}0&\hat{1}^{-{\dagger}}\\ \mp\hat{1}&0\end{pmatrix}$ $\in{\mathsf{U}}^{\mp}(1)\subseteq{\mathsf{U}}^{\mp}(q)$
@@
Let us very briefly review the transcription of Pauli groups to $R$-modules [31]. The group of all finitely supported Pauli operators on ${\mathbb{Z}}^{\mathsf{d}}$ with $q$ qudits ${\mathbb{C}}^{p}$ per site is nonabelian, but its commutator subgroup consists of $p$-th root of unity. Since the single-site Pauli group abelianizes to ${\mathbb{F}}_{p}^{2q}$, the abelianization of the Pauli group on the infinite lattice is the additive group of $R^{2q}$. The translation group action gives an $R$-module structure to $R^{2q}$ that matches the usual ring action on the free $R$-module $R^{2q}$. It is our convention that upper half block of a Laurent polynomial column matrix, (i.e., $R^{q}\oplus 0\subseteq R^{q}\oplus R^{q}$), corresponds to the “$X$”-part, operator factors $X=\sum_{j\in{\mathbb{F}}_{p}}\ket{j+1}\bra{j}$ | and its powers, and the lower half block $0\oplus R^{q}$ to the “$Z$”-part, operator factors $Z=\sum_{j\in{\mathbb{F}}_{p}}e^{2\pi\mathrm{i}j/p}\ket{j}\bra{j}$ |. For example,
@@
If a ${\mathbb{C}}$-algebra automorphism $\beta$ maps a Pauli operator to a Pauli operator, we obtain a map $U$ from a Laurent polynomial column matrix to another. If $\beta$ is translation invariant, then the induced map $U$ must be $R$-linear so we have an $R$-module automorphism on $R^{2q}$. The size of a Pauli operator under $\beta$ can only be finite, and since the translation invariant map $\beta$ is determined by finitely many images, the locality blowup is uniformly bounded across the whole lattice ${\mathbb{Z}}^{\mathsf{d}}$. Therefore, $\beta$ is a Clifford QCA. Since any ${\mathbb{C}}$-algebra automorphism must preserve commutation relation, the induced $R$-automorphism $U$ must preserve $\lambda$. We hence obtain a map from Clifford QCA to $\lambda$-unitaries. Introducing auxiliary qudits in the lattice on which a QCA acts by identity amounts to our stabilization.
@@
It suffices to consider $U^{\prime}=I$ because we may multiply everything by $(U^{\prime})^{-1}$. By definition of ${\mathsf{U}}^{\mp}(q)$, we must have $U=\begin{pmatrix}I&Y\\ 0&I\end{pmatrix}$ where $Y^{\dagger}\mp Y=0$. If $U\in{\widetilde{\mathsf{U}}}^{\mp}(q)$, then $Y=\theta\pm\theta^{\dagger}$ for some $\theta$. ∎
@@
Suppose $L\subset Q\oplus Q^{*}$ is a lagrangian of a trivial quadratic $\mp$-form $\eta$ on $Q\oplus Q^{*}$ where $Q$ is a finitely generated projective $R$-module. Then, there exists an $\eta^{\mp}$-unitary $U$ on $Q\oplus Q^{*}$ such that $U(Q\oplus 0)=L$ and such $U$ is unique up to ${\widetilde{\mathsf{EU}}}^{\mp}(Q)$. Similarly, if $L$ is a lagrangian with respect to the trivial hermitian form $\lambda^{\mp}$ on $Q\oplus Q^{*}$, then there exists a $\lambda^{\mp}$-unitary $U$ such that $U(Q\oplus 0)=L$, unique up to ${\mathsf{EU}}^{\mp}(Q)$.